OpenStudy (anonymous):

2x + 5y = −3 −15x − 5y = −10

OpenStudy (rmrjr22):

find x and y?

OpenStudy (anonymous):

Find x & y?

OpenStudy (anonymous):

yup...but my stupid teacher needs all the work xc

OpenStudy (rmrjr22):

cancel out a variable x or y

OpenStudy (rmrjr22):

in this case y is all ready to go

OpenStudy (rmrjr22):

so ur left with x to solve

OpenStudy (rmrjr22):

add both equations together

OpenStudy (anonymous):



OpenStudy (anonymous):

hehe thanks guys <3

OpenStudy (rmrjr22):

do you get it?

OpenStudy (anonymous):

what do you get for x? @crilea93

OpenStudy (anonymous):

yes thank you c: both of you.. and doesnt x=1?

OpenStudy (rmrjr22):

yes

OpenStudy (anonymous):



OpenStudy (anonymous):

2x + 5y = -3 2(1) + 5y = -3 2 + 5y = -3 -2 -2 ---- ---- 5y = -5 Divide by 5 5y = -5 -- --- 5 5 y = ?

OpenStudy (anonymous):

Then you will have: x = 1 & y = __ (x, y) (1, __)

OpenStudy (anonymous):

2x + 5y = −3 −15x − 5y = −10 -13x=-13 x=1 2x + 5y = -3 2(1) + 5y = -3 2 + 5y = -3 -2 -2 ---- ---- 5y = -5 y=-1?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

x = 1 & y = -1

OpenStudy (anonymous):

soo... (1,-1)

OpenStudy (anonymous):

Thank you c:

OpenStudy (rmrjr22):

bam!

OpenStudy (anonymous):

(x, y) (1, -1) yup

OpenStudy (anonymous):

yw :)

OpenStudy (anonymous):

@crilea93 {Best Response} You give a medal to whomever you think helped you with you question :)